OpenStudy (anonymous):

find the equation of the tangent to the curve y=(e^x/3)/x at the point (3,1/3e)

zepdrix (zepdrix):

Ok soooo, to find the tangent line, we first need the derivative of the function. Do you know how to get the derivative of that function? :)

OpenStudy (anonymous):

not for the e^x function ..

zepdrix (zepdrix):

so it looks like you have functions being divided :D so we'll need to apply the quotient rule yes? (d/dx)e^x = e^x, remember how to setup the quotient rule? :O oh quick question... is the exponential part... \[e^{x/3}\]or is it..\[\frac{ e^x }{ 3 }\]

OpenStudy (anonymous):

its \[e ^{x/3}/x\]

zepdrix (zepdrix):

oh ok.. lets look at the derivative of the exponential term before starting this one :D \[(d/dx)e^{x/3}=e^{x/3}(d/dx)(\frac{ x }{ 3 })=e^{x/3}(\frac{ 1 }{ 3 })\] ok confused about anything above?

OpenStudy (anonymous):

yeah i actually am ..sorry.. what about if it was \[e ^{x/3}\]

zepdrix (zepdrix):

that's what we're looking at right now, im ignoring the bottom x for a moment.

zepdrix (zepdrix):